Circular No. 8406

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004el, 2004em, 2004en, 2004eo, 2004ep Five apparent supernovae have been discovered on unfiltered CCD im-

ages: 2004el in MCG +09-25-4 by R. Arbour (cf. IAUC 8376), 2004em in IC 1303 by M. Armstrong (cf. IAUC 8399), 2004en by T. Boles (cf. IAUC 8405), 2004eo in NGC 6928 by K. Itagaki (cf. IAUC 8377; via S. Nakano), and 2004ep in IC 2152 by M. Moore and W. Li (LOSS/KAIT; cf. IAUC 8401).

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004el Sept. 14.870 14^{h}59^{m}51.^{s}79 +54^{o}37′06′′.1 16.8 9′′.9 W, 2′′.7 S
2004em Sept. 14.968 19 31 31.11 +35 52 15.7 17.5 10′′.4 E, 20′′.3 S
2004en Sept. 16.054 2 36 26.53 +35 59 25.5 18.2 8′′.9 E, 1′′.0 N
2004eo Sept. 17.500 20 32 54.19 + 9 55 42.7 17.8 59′′.1 E, 6′′.5 N
2004ep Sept. 22.52 5 57 53.66 −23 10 58.3 18.6 3′′.8 E, 7′′.2 S

Approximate unfiltered CCD magnitudes (from the respective discoverers unless otherwise noted): SN 2004el, Mar. 30 UT, [19.5 (Boles); 6.92, [17.5; Sept. 14.943, 17.2 (Boles); 15.822, 16.8. SN 2004em, 1988 June 13, [22.5 (blue Palomar Sky Survey); 1992 Aug. 31, [20.8 (red Palomar Sky Survey); 2004 July 28, [19.0; Sept. 15.831, 17.5. SN 2004en, 1986 Nov. 27, [21.0 (red Digitized Sky Survey); 1987 Oct. 1, [20.5 (blue Digitized Sky Survey); 2003 Nov. 27, [19.5; Dec. 28, [19.5; 2004 Sept. 16.966, 18.2. SN 2004eo, 2002 July 30, [18.5; 2004 Sept. 8, [18.5; 12, [18.5; 18.647, 17.6 (poor weather conditions). SN 2004ep, Mar. 5.20, [19.5; Aug. 16.54, 18.7. An image of SN 2004el taken by D. Briggs (near Portsmouth, England) on Sept. 14.9 yielded position end figures 51^{s}.84, 05′′.7. SN 2004el was also not present on the (Digitized Sky Survey) Palomar plates taken in 1991 (limiting blue mag 20.1) and 1994 (limiting red mag 19.8).

M. Modjaz, P. Challis and R. Kirshner, Harvard-Smithsonian Center for Astrophysics; and T. Matheson, National Optical Astronomy Observatory, report that a spectrum (range 350–740 nm) of SN 2004el, obtained by P. Berlind on Sept. 17.12 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), reveals it to be a supernova of type II. The spectrum consists of a blue continuum with P-Cyg lines of Hα and Hβ. Adopting the Lyon-Meudon Extragalactic Database recession velocity of 7885 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 9900 km/s.

2004 September 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green